A conducting and closed container of volume 300 litres contain an ideal gas at a high pressure Pinitial . Now using a pump , the gas it taken out at a constant rate 5 litre per sec . find the time in which the pressure inside the container decreases to Pinitial/100 .

1
Expert's answer
2022-04-11T12:02:51-0400

"p_1V_1=p_2V_2,"

"\frac{V_1}{V_2}=100,"

"V_2=\frac{V_1}{100}=3~l,"

"\u2206V=V_1-V_2=297~l,"

"t=\frac{∆V}v=1~min."